A coil with magnetic moment 1.42 A⋅m2 is oriented initially with its magnetic moment antiparallel to a uniform magnetic field of magnitude 0.810 T .

What is the change in potential energy of the coil when it is rotated 180 degrees, so that its magnetic moment is parallel to the field?

Express your answer in joules.
Answer #1

U -= mu*B*cos phu

U1 = mu*B*cos 0 = mu*B

U2 = mu*B*cos 180 = -mu*B

dU = U2 - U1 = -2*mu*B

dU = -2.0*1.42*0.810 = -2.3004 J